OpenStudy (anonymous):

NEED HELP! WILL FAN AND MEDAL! Solve system by elimination. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 I know the answer is (1,1,0), but I don't know the steps to get there.

OpenStudy (anonymous):

@kiamousekia @okseeyah @Tyler.Evan2 Do you guys know this?

OpenStudy (anonymous):

Im sorry, IDK

OpenStudy (anonymous):

What grade is this?

OpenStudy (anonymous):

Alegbra 2, 10th grade. @Tyler.Evan2

OpenStudy (anonymous):

Oh, LOL im in 7th advanced math.

OpenStudy (word2):

what is the question ?

OpenStudy (anonymous):

the top one ^^^^ @word2

OpenStudy (word2):

hold on

OpenStudy (word2):

okay , for the first one i solved for x by simplifying both sides of the equation , then isolating the variable i got \[x=y+3z/2\]

OpenStudy (word2):

for the second one i did the same thing again and got x=-y/2+z/2+3/2

OpenStudy (word2):

and the last one i got x = -3y/2 - 3z/2 =5/2